# BELT: Old-School Backdoor Attacks can Evade the State-of-the-Art Defense with Backdoor Exclusivity Lifting

### v1

@@ 6.3.1 Neural Cleanse
We evaluate all BadNet models trained in Section 6.2 using Neural Cleanse [1]. We conduct 100 optimization periods through a reverse engineering process to generate reverse triggers for each class of the evaluated models. The neural cleanup process is repeated 10 times for each model to obtain an average anomaly metric, as illustrated in  Figure 5. The results reveal that the anomaly index of the BELT model for all settings remains below the threshold of 2. This suggests that Neural Cleanse fails to detect the backdoor in the BELT model, whether it is a ${}^{+}$ or a ${}^{++}$. The reason for this failure lies in Neural Cleanse’s inability to accurately reverse engineer the fuzzy trigger through the reverse engineering process.  Figure 6 provides an example set of reverse triggers for the target tag on the CIFAR-10 task, where the first column represents the original trigger for the backdoor. The benign model, lacking a backdoor, presents the reverse triggers as natural features of the target class, while the BadNet model, with a exclusivity of 0%, allows fuzzy triggers to be easily reverse-engineered. However, for ${}^{+}$ and ${}^{++}$, the reverse triggers of the target class closely resemble those of the benign model. This is attributed to BELT significantly reducing the presence of fuzzy triggers by strengthening the backdoor exclusivity, making it challenging for Neural Cleanse to identify fuzzy triggers within the narrowed exclusivity boundary during the reverse engineering process.
@@ 6.3.2 ABS
It is observed that ABS can effortlessly detect the backdoor in the regular Badnet model, as its REASR exceeds the threshold of 0.88 and approaches 1.0. However, ABS fails to effectively detect ${}^{+}$ and ${}^{++}$, where their REASR values are close to those of the benign model. This is not surprising; similar to Neural Cleanse, ABS performs reverse engineering on compromised neurons to obtain corresponding reverse triggers. The strong exclusivity of the backdoor hinders this process, preventing it from acquiring a reverse trigger capable of activating the backdoor, resulting in REASR values far below the threshold.
@@ 6.3.3 MNTD
${}^{++}$ 62.44 95.44 7.55
For the evaluation of Meta-Neural Trojan Detection (MNTD) [5], we adhere to the training settings of MNTD initially, employing transfer learning to construct a meta-classifier. This meta-classifier is trained on the CIFAR-10 dataset, comprising 2048 benign models and 2048 backdoor models within a shadow model set. During the evaluation phase, we train 256 benign test models (referred to as ”Benign”) and 256 backdoor test models (referred to as ”Backdoored”). Additionally, we train an additional 256 backdoor models in the data outsourcing scenario (referred to as ${}^{+}$”) and another 256 backdoor models in the model outsourcing scenario (referred to as ${}^{++}$”).
It is crucial to highlight that both ${}^{+}$” and ${}^{++}$” maintain identical training settings as ”Backdoored,” encompassing model architecture, backdoor configurations, and training hyperparameters. This ensures that any impact on detection performance stems solely from the actions of the backdoor itself, rather than changes in hyperparameters [44]. In Table III, the average performance of the three test model sets is presented. MNTD demonstrates commendable performance on the test set containing ordinary backdoored models (”Backdoored”), achieving an AUC of 95.63%. However, the high-exclusivity backdoor models effectively resist MNTD detection, exhibiting AUCs of 75.81% and 77.61% for the test sets based on ${}^{+}$” and ${}^{++}$,” respectively. The primary cause lies in the fact that backdoors with stronger exclusivity exhibit reduced responsiveness to optimized queries compared to those with weaker exclusivity. Consequently, meta-classifiers are inclined to predict lower scores for strongly exclusive backdoor models, posing challenges in distinguishing them significantly from benign models and leading to diminished detection performance.
@@ 6.3.5 STRIP
Figure 8 displays the entropy distribution results of benign samples and trigger-carrying samples on a regular Badnet and a exclusivity-enhanced Badnet as evaluated by STRIP [2]. The assessed models utilize the ResNet18 architecture and are trained on the CIFAR-10 dataset. The distribution is evaluated on 1000 trigger images and 1000 clean images. For the regular Badnet model, STRIP effectively distinguishes samples carrying triggers from clean samples because there is a noticeable gap in the entropy value distribution between trigger samples and clean samples. However, for ${}^{+}$ and ${}^{++}$, their entropy distributions are similar to the benign model, with a significant overlap, rendering STRIP ineffective. The primary reason for this phenomenon is that the high exclusivity of the backdoor disrupts the strong hijacking effect of the trigger under different perturbations. As a result, the perturbed trigger, when overlaid with benign samples, is no longer a fuzzy trigger capable of activating the backdoor.
@@ 6.3.6 SentiNet
We partition the clean data into intervals based on a fixed average confidence step and select the two data points with the highest Foolde within each interval as decision points for fitting the decision boundary. Figure Figure 9 illustrates the detection performance of SentiNet [4]. For a regular Badnet, the trigger on poisoned data significantly influences the model output, resulting in high average confidence and Foolde within each interval. This leads to an effective decision boundary that distinguishes poisoned data from clean data. However, the detection performance of SentiNet is not satisfactory for ${}^{+}$ and ${}^{++}$. Poisoned data and clean data are confounded, causing the decision boundary to be ineffective in distinguishing them. One possible reason is that GradCAM cannot ensure the capture of a complete trigger (see the last column of Figure Figure 1). For highly exclusive backdoors, such incomplete triggers do not elicit a response from the backdoor, escaping detection by SentiNet.
@@ 7 Discussion
  • Comparison Group C: Without dynamic weight. We use this baseline to validate the effectiveness of dynamic weights. Here, we remove the dynamic weight mechanism from Equation 7 and replace it with a small static value (0.1). The table shows a significant performance drop on both ${}^{+}$ and ${}^{++}$, confirming that dynamic weights help bring the perturbation closer to the trigger upper bound during the optimization process.
@@
We use these two approaches to build two new cover datasets, and each is used to train a Badnet backdoored model. The models use a ResNet18 architecture and the CIFAR-10 dataset. The final backdoor exclusivity achieved by these models is 8.10% and 14.37%, respectively. In comparison, the cover triggers constructed based on the masking approach show the most significant improvement in exclusivity, reaching 16.46% under the same settings. The improvement in exclusivity with cover triggers constructed through random sampling is limited, mainly because it is challenging to construct cover triggers in a controlled manner using this approach. In contrast, cover triggers constructed by adding noise and the masking approach can be controlled by adjusting the noise intensity and masking ratio. However, constructing cover triggers based on adding noise is challenging to determine the appropriate noise intensity. We opt for the masking-based approach, using a default mask rate of 20%, which achieves good exclusivity enhancement in all settings.
(d) Limitations and Future Works. In this study, we have taken a step towards evaluating and fortifying backdoor exclusivity in the context of data outsourcing and model outsourcing scenarios. To precisely quantify backdoor exclusivity is computationally challenging. We instead design an optimization-based algorithm to estimate the exclusivity score. Our algorithm incorporates optimization directions and dynamic weights to evade the potential local minima during the estimation. The experimental results show the estimated exclusivity score does reflect well the stealthiness of different backdoor attacks. Considering the importance of the exclusivity score in characterizing the backdoor stealthiness, it would be meaningful for future works to devise more accurate estimation of backdoor exclusivity via, .e.g, more advanced numeric techniques.
@@ 8 Conclusion
In this work, we propose and investigate a novel characteristic of backdoor attacks called backdoor exclusivity. Based on our proposed approximation algorithm to quantify exclusivity, we conduct a comprehensive analysis of the intricate relation between backdoor exclusivity and the existence of many fuzzy triggers in the classical backdoor attacks, which are intensively exploited by the state-of-the-art defenses to detect and recover the triggers. To enhance the capability of existing backdoor attacks to evade the state-of-the-art defenses, we introduce Backdoor Exclusivity LifTing (BELT) which suppresses the association between the backdoor and the potential fuzzy triggers and can be seamless combined with almost any existing backdoor attacks. We thoroughly validate the effectiveness of BELT on four existing backdoor attacks which would otherwise be detected by the state-of-the-art defenses. After applying the BELT technique, the backdoored models exhibit high evasiveness against six state-of-the-art backdoor defenses. This is primarily attributed to BELT refining the activation conditions of the backdoor, disrupting the necessary assumptions of these defense mechanisms, and causing them to struggle in obtaining an effective fuzzy trigger in the presence of a strongly exclusive backdoor.

### v2

@@ 6.3.1 Neural Cleanse
We evaluate all BadNet models trained in Section 6.2 using Neural Cleanse [2]. We conduct 100 optimization periods through a reverse engineering process to generate reverse triggers for each class of the evaluated models. The neural cleanup process is repeated 10 times for each model to obtain an average anomaly metric, as illustrated in  Figure 5. The results reveal that the anomaly index of the BELT model for all settings remains below the threshold of 2. This suggests that Neural Cleanse fails to detect the backdoor in the BELT model, whether it is a BadNet+ or a BadNet++. The reason for this failure lies in Neural Cleanse’s inability to accurately reverse engineer the fuzzy trigger through the reverse engineering process.  Figure 6 provides an example set of reverse triggers for the target tag on the CIFAR-10 task, where the first column represents the original trigger for the backdoor. The benign model, lacking a backdoor, presents the reverse triggers as natural features of the target class, while the BadNet model, with a exclusivity of 0%, allows fuzzy triggers to be easily reverse-engineered. However, for BadNet+ and BadNet++, the reverse triggers of the target class closely resemble those of the benign model. This is attributed to BELT significantly reducing the presence of fuzzy triggers by strengthening the backdoor exclusivity, making it challenging for Neural Cleanse to identify fuzzy triggers within the narrowed exclusivity boundary during the reverse engineering process.
@@ 6.3.2 ABS
It is observed that ABS can effortlessly detect the backdoor in the regular Badnet model, as its REASR exceeds the threshold of 0.88 and approaches 1.0. However, ABS fails to effectively detect Badnet+ and Badnet++, where their REASR values are close to those of the benign model. This is not surprising; similar to Neural Cleanse, ABS performs reverse engineering on compromised neurons to obtain corresponding reverse triggers. The strong exclusivity of the backdoor hinders this process, preventing it from acquiring a reverse trigger capable of activating the backdoor, resulting in REASR values far below the threshold.
@@ 6.3.3 MNTD
Backdoored++ 62.44 95.44 7.55
For the evaluation of Meta-Neural Trojan Detection (MNTD) [5], we adhere to the training settings of MNTD initially, employing transfer learning to construct a meta-classifier. This meta-classifier is trained on the CIFAR-10 dataset, comprising 2048 benign models and 2048 backdoor models within a shadow model set. During the evaluation phase, we train 256 benign test models (referred to as ”Benign”) and 256 backdoor test models (referred to as ”Backdoored”). Additionally, we train an additional 256 backdoor models in the data outsourcing scenario (referred to as ”Backdoored+”) and another 256 backdoor models in the model outsourcing scenario (referred to as ”Backdoored++”).
It is crucial to highlight that both ”Backdoored+” and ”Backdoored++” maintain identical training settings as ”Backdoored,” encompassing model architecture, backdoor configurations, and training hyperparameters. This ensures that any impact on detection performance stems solely from the actions of the backdoor itself, rather than changes in hyperparameters [45]. In Table III, the average performance of the three test model sets is presented. MNTD demonstrates commendable performance on the test set containing ordinary backdoored models (”Backdoored”), achieving an AUC of 95.63%. However, the high-exclusivity backdoor models effectively resist MNTD detection, exhibiting AUCs of 75.81% and 77.61% for the test sets based on ”Backdoored+” and ”Backdoored++,” respectively. The primary cause lies in the fact that backdoors with stronger exclusivity exhibit reduced responsiveness to optimized queries compared to those with weaker exclusivity. Consequently, meta-classifiers are inclined to predict lower scores for strongly exclusive backdoor models, posing challenges in distinguishing them significantly from benign models and leading to diminished detection performance.
@@ 6.3.5 STRIP
Figure 8 displays the entropy distribution results of benign samples and trigger-carrying samples on a regular Badnet and a exclusivity-enhanced Badnet as evaluated by STRIP [3]. The assessed models utilize the ResNet18 architecture and are trained on the CIFAR-10 dataset. The distribution is evaluated on 1000 trigger images and 1000 clean images. For the regular Badnet model, STRIP effectively distinguishes samples carrying triggers from clean samples because there is a noticeable gap in the entropy value distribution between trigger samples and clean samples. However, for Badnet+ and Badnet++, their entropy distributions are similar to the benign model, with a significant overlap, rendering STRIP ineffective. The primary reason for this phenomenon is that the high exclusivity of the backdoor disrupts the strong hijacking effect of the trigger under different perturbations. As a result, the perturbed trigger, when overlaid with benign samples, is no longer a fuzzy trigger capable of activating the backdoor.
@@ 6.3.6 SentiNet
We partition the clean data into intervals based on a fixed average confidence step and select the two data points with the highest Foolde within each interval as decision points for fitting the decision boundary. Figure Figure 9 illustrates the detection performance of SentiNet [7]. For a regular Badnet, the trigger on poisoned data significantly influences the model output, resulting in high average confidence and Foolde within each interval. This leads to an effective decision boundary that distinguishes poisoned data from clean data. However, the detection performance of SentiNet is not satisfactory for Badnet+ and Badnet++. Poisoned data and clean data are confounded, causing the decision boundary to be ineffective in distinguishing them. One possible reason is that GradCAM cannot ensure the capture of a complete trigger (see the last column of Figure Figure 1). For highly exclusive backdoors, such incomplete triggers do not elicit a response from the backdoor, escaping detection by SentiNet.
6.3.7 Fine-Pruning
Fine-Pruning [34] is one of the earliest defenses based on parameter updating, which is motivated from the following observation: the inputs with triggers usually activate neurons that are dormant in the presence of clean inputs. Therefore, Fine-Pruning attempts to erase the backdoor by pruning neurons which are scarcely activated for the clean inputs. To reduce the impact on CDA, Fine-Pruning further applies fine-tuning after pruning to restore the model performance.
To verify the impact of updating model parameters on BELT, we apply Fine-Pruning (with the default configuration in its official implementation [1]) on BadNet+ and BadNet++, with the CIFAR10 dataset and the ResNet18 architecture. As shown in Figure 10, Fine-Pruning fails to effectively remove the backdoor enhanced by BELT, since ASR remains higher than CDA at various pruning rates. Only when the pruning rate reaches 90% does ASR begin to decrease, while CDA has already suffered an unacceptable loss.
Figure 10: Performance of Fine-Pruning on backdoored models after BELT.
@@ 7 Discussion
  • Comparison Group C: Without dynamic weight. We use this baseline to validate the effectiveness of dynamic weights. Here, we remove the dynamic weight mechanism from Equation 7 and replace it with a small static value (0.1). The table shows a significant performance drop on both BadNet+ and BadNet++, confirming that dynamic weights help bring the perturbation closer to the trigger upper bound during the optimization process.
@@
We use these two approaches to build two new cover datasets, and each is used to train a Badnet backdoored model. The models use a ResNet18 architecture and the CIFAR-10 dataset. The final backdoor exclusivity achieved by these models is 58.43% and 80.05%, respectively. In comparison, the cover triggers constructed based on the masking approach show the most significant improvement in exclusivity, reaching 84.55% under the same settings. The improvement in exclusivity with cover triggers constructed through random sampling is limited, mainly because it is challenging to construct cover triggers in a controlled manner using this approach. In contrast, cover triggers constructed by adding noise and the masking approach can be controlled by adjusting the noise intensity and masking ratio. However, constructing cover triggers based on adding noise is challenging to determine the appropriate noise intensity. We opt for the masking-based approach, using a default mask rate of 20%, which achieves good exclusivity enhancement in all settings.
(d) Adaptive Defense. In this part, we analyze the potential adaptive defense against BELT. As the effectiveness of BELT relies on the exclusivity of backdoors, only sufficiently precise triggers have the ability to activate the backdoor. In other words, the backdoor exclusivity prevents the backdoor from responding to fuzzy triggers. Therefore, an adaptive defender may attempt to add noises to the input data to make the potential triggers fuzzy, thereby circumventing the backdoor effect. Below, we further verify the resilience of BELT to Gaussian perturbations. As shown in Figure 13, for BadNet+ under the CIFAR-10 and ResNet18 settings, CDA/ASR decreases from 94.85%/99.98% to 69.79%/90.54% as the standard deviation of the noise increases from 00 to $0.05$. BadNet++ exhibits the same trend as BadNet+, with ASR consistently outperforming CDA as the noise intensity increases. This indicates that BELT achieves balance between backdoor exclusivity and perturbation resistance.
Figure 13: Performance of BELT in Gaussian noise with different standard deviations.
(e) Limitations and Future Works. In this study, we have taken a step towards evaluating and fortifying backdoor exclusivity in the context of data outsourcing and model outsourcing scenarios. To precisely quantify backdoor exclusivity is computationally challenging. We instead design an optimization-based algorithm to estimate the exclusivity score. Our algorithm incorporates optimization directions and dynamic weights to evade the potential local minima during the estimation. The experimental results show the estimated exclusivity score does reflect well the stealthiness of different backdoor attacks. Considering the importance of the exclusivity score in characterizing the backdoor stealthiness, it would be meaningful for future works to devise more accurate estimation of backdoor exclusivity via, .e.g, more advanced numeric techniques.
@@ 8 Conclusion
In this work, we propose and investigate a novel characteristic of backdoor attacks called backdoor exclusivity. Based on our proposed approximation algorithm to quantify exclusivity, we conduct a comprehensive analysis of the intricate relation between backdoor exclusivity and the existence of many fuzzy triggers in the classical backdoor attacks, which are intensively exploited by the state-of-the-art defenses to detect and recover the triggers. To enhance the capability of existing backdoor attacks to evade the state-of-the-art defenses, we introduce Backdoor Exclusivity LifTing (BELT) which suppresses the association between the backdoor and the potential fuzzy triggers and can be seamless combined with almost any existing backdoor attacks. We thoroughly validate the effectiveness of BELT on four existing backdoor attacks which would otherwise be detected by the state-of-the-art defenses. After applying the BELT technique, the backdoored models exhibit high evasiveness against seven state-of-the-art backdoor defenses. This is primarily attributed to BELT refining the activation conditions of the backdoor, disrupting the necessary assumptions of these defense mechanisms, and causing them to struggle in obtaining an effective fuzzy trigger in the presence of a strongly exclusive backdoor.